OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -51 ? -126 -123 -120 -117

OpenStudy (anonymous):

\[-51=a_{17}=a_1+16d=13+16d\]\[-51=13+16d\] \[-64=16d\] \[d=-4\] \[a_{35}=a_1+34d=13+(-4)34=13-136=-123\]